# Auditing a Dutch Public Sector Risk Profiling Algorithm Using an Unsupervised Bias Detection Tool

### v2

@@ Abstract
Algorithms are increasingly used to automate or aid human decisions, yet recent research shows that these algorithms may exhibit bias across legally protected demographic groups. However, data on these groups may be unavailable to organizations or external auditors due to privacy legislation. This paper studies bias detection using an unsupervised clustering tool when data on demographic groups are unavailable. We collaborate with the Dutch Executive Agency for Education to audit an algorithm that was used to assign risk scores to college students at the national level in the Netherlands between 2012-2023. Our audit covers more than 250,000 students from the whole country. The unsupervised clustering tool highlights known disparities between students with a non-European migration background and Dutch origin. Our contributions are three-fold: (1) we assess bias in a real-world, large-scale and high-stakes decision-making process by a governmental organization; (2) we use simulation studies to highlight potential pitfalls of using the unsupervised clustering tool to detect true bias when demographic group data are unavailable and provide recommendations for valid inferences; (3) we provide the unsupervised clustering tool in an open-source library. Our work serves as a starting point for a deliberative assessment by human experts to evaluate potential discrimination in algorithmic-supported decision-making processes.
Keywords fairness  $\cdot$ auditing  $\cdot$ bias  $\cdot$ clustering  $\cdot$ indirect discrimination
@@ 1 Introduction
The use of algorithmic support for decision-making has grown in recent years [62], influencing decisions such as who gets a loan, who makes it to the next recruitment stage for a job, and who is flagged for fraud investigation [61, 57, 65]. As algorithmic systems are adopted, their negative impacts have also become more clear [9, 68, 45, 7, 57]. These systems may replicate the bias we see in our society, leading to potentially discriminatory effects [63]. In this study, we outline a methodology to detect such biases in algorithms without data on demographic groups using an unsupervised bias detection algorithm, evaluated on a Dutch public sector risk profiling algorithm.
@@
In the period 2012-2023, a simple rule-based risk profiling algorithm from the Dutch Executive Agency for Education (DUO) contributed to indirect discrimination in a control process to check whether students were (un)duly allocated a college grant [6]. A rule-based algorithm, consisting of three apparently neutral characteristics of students (type of education, age, and distance to parents), was used to assign higher risk scores to students enrolled in vocational education, younger students, and students who lived near their parental address. The seemingly impartial risk assessment led to unequal treatment of students with a non-European migration background, remaining unnoticed for over 10 years and affecting more than 350,000 students [^1] Between 2012-2023, over 350,000 students were assigned a risk score by the algorithm. Our analysis focuses on the student population in two years: 2014 and 2019, which consists of a total of 298,882 students. [5].
@@
Given this, in 2024, the Dutch Minister for Education, Culture and Science apologized on behalf of the Dutch government for indirect discrimination experienced by students with a non-European migration background and announced a EUR 61 million compensation plan [5]. This example of a new discriminatory algorithm scandal in the Netherlands [35] highlights that not only complex statistical methods, like machine learning and deep learning [16], but also simple rule-based algorithms pose significant risks of embedding discrimination in algorithm-supported decision-making processes. In a recent interview, the chair of the Dutch Data Protection Authority warned that discriminatory rule-based algorithms are “around every corner” in executive branches of the Dutch government [60].
This type of risk profiling is prohibited by the European Convention of Human Rights (ECHR), the Charter of Fundamental Rights of the European Union (EU) [3], the EU General Data Protection Regulation (GDPR) [4] and the EU AI Act [29], as well as the national Public Administration Law [2]. Limited institutional enforcement capabilities, along with a lack of awareness, have led to inadequate oversight of algorithmic-supported decision-making processes [48].
@@
To combat bias in algorithmic systems, the academic community has presented a multitude of ways to define, detect, and mitigate different notions of bias at different points along the algorithmic pipeline [47, 18, 11, 69]. Most bias mitigation methods require data on the demographic groups to prevent discriminatory bias. These demographic groups are typically protected attributes under discrimination law, such as race, gender and age [^2] The exact protected grounds are enshrined in European non-discrimination law, such as the Treaty on the Functioning of the European Union (TFEU), the Charter of Fundamental Rights of the EU, and specific directives like the Racial Equality Directive (2000/43/EC) and the Employment Equality Directive (2000/78/EC). The main protected attributes include: race and ethnicity, gender and sex, religion or belief, disability, age, sexual orientation, nationality, and marital or family status. Throughout, we will refer to these attributes as demographic groups. Note that some provisions do not mention any particular grounds of discrimination, such as Art. 20 of the EU Charter. Instead, using openly formulated clauses such as ”everyone is equal before the law”. [10]. However, data on demographic groups are often not available to organizations or external algorithm auditors [8, 15]. Upon request, public sector organizations can occasionally obtain aggregated statistics on demographic groups through official institutions, such as the national office of statistics, which have access to statistical population registers. Private organizations cannot request these data and face ethical, legal, and practical challenges when collecting data on demographic groups customers belong to  [15, 30, 29]. Unsupervised learning can aid in this challenge by detecting groups in the data for which a given notion of equal treatment is violated, avoiding the need for ex-ante specification of demographic or legally protected groups [49].
@@
In this paper, we use an unsupervised bias detection tool to audit the risk profiling algorithm from DUO for bias. For this audit, DUO provided data processed by their algorithm that is not publicly accessible, and Statistics Netherlands provided the aggregate data on country of birth and country of origin [52]. These data enable us to verify whether the clusters identified by the unsupervised bias detection tool, which represent potentially unfairly treated groups, correspond to disadvantaged demographic groups. This allows for a comparison between an unsupervised learning approach and an approach using demographic data. We emphasize that the clustering procedure itself does not rely on the available protected demographic data: we only use the aggregated demographic data to illustrate, post hoc, that the flagged clusters align with negatively affected groups. This analysis is therefore a validation of the clustering methodology, not a replacement for it.
In our context, the bias metric [^3] Throughout our paper, we say bias metric but this can be understood as a fairness metric as commonly used in the algorithmic fairness literature. The fairness metric reflected by our bias metric definition is demographic parity [22]. that we use requires that the probability of a student with a Dutch or European migration background being classified as ‘high risk’ by the risk profiling algorithm is equal to the probability of a student with a non-European migration background being classified as ‘high risk’. We justify the use of this metric since it assesses the adverse effect of the risk profiling algorithm on different demographic groups: a crucial aspect for investigating indirect discrimination under European non-discrimination law.
@@
We find that the cluster with the highest bias returned by the unsupervised bias detection tool is characterized by students who are more often enrolled in vocational education, live close to their parental address, and are relatively more often 15-18 or 25-50 years old. This aligns with demographic groups found to be negatively affected by the risk profiling algorithm, based on aggregated statistics regarding students’ non-European migration background.
Our paper makes three contributions:
  1. 1.
    Large scale audit of indirect discrimination: The DUO audit presents a high-quality and socially relevant dataset, providing insights into the assessment of indirect discrimination in algorithm-supported decision-making. This study thereby contributes to the advancement of public knowledge building and establishment of standards for auditing algorithmic systems on indirect discrimination.
  2. 2.
    Advancing methodology for unsupervised bias detection: We build on existing clustering algorithms and strengthen their methodology. Through using a simulation study, we examine conditions under which the tool works best and highlight key factors to consider for reliable inferences when applying it. Our methodological improvements substantially reduce the false positive rate of the algorithm.
  3. 3.
    Open-source tool: We have published the unsupervised bias detection method as an open-source tool in the form of a Python package, as well as an online web application. This allows a non-technical audience to easily use the tool on their own datasets to detect potential bias.
@@
Overall, our paper showcases the potential of unsupervised learning to aid in auditing algorithmic-supported decision-making processes via an application to a real-world use case. We argue that unsupervised bias detection serves as a starting point for a deliberative assessment by human experts to evaluate potential discrimination and unfairness in the algorithmic-supported decision-making process under review when demographic groups are unavailable.
The rest of the paper is outlined as follows. We start by reviewing related literature in Section 2. We present the methodology behind the unsupervised bias detection tool in Section 3. Then, in Section 4, we describe the application of the unsupervised bias detection tool to DUO’s risk profiling algorithm. We follow this by providing simulation studies in Section 5 which help to understand potential pitfalls when using the tool. We finish with a discussion on the limitations and conclusions of our work in Sections 7 and 8.
@@ 2 Literature Review
We review a selection of the literature on bias detection and non-discrimination law, structuring the review around three topics: (1) challenges in detecting and establishing indirect discrimination, (2) bias detection methods with and without access to demographic data, and (3) auditing algorithmic systems on bias.
@@ 2.1 Challenges in Detecting and Establishing Discrimination
Under European non-discrimination law, DUO’s control process is classified as indirect discrimination as an apparently neutral practice disproportionately disadvantaged students with a non-European migration background [^4] Supra note 2 [54]. In algorithms, discrimination often arises from the proxy and correlation challenge [31]. Proxies refer to characteristics that are correlated with a demographic group, leading to disadvantaged outcomes for this group. For example, the Court of Justice of the European Union (CJEU) found that using part-time employment as a basis for differentiation in the context of occupational pension schemes constituted indirect discrimination, as over 80% of the part-time workers in Spain in 2012 were women [55]. In some cases, courts may accept that differential treatment of a demographic group is acceptable. Under EU law, indirect discrimination can be justified if it is demonstrated that a legitimate aim is being pursued, and the means of achieving that aim are appropriate and necessary [24, 25, 26, 27]. Our paper highlights how unsupervised methods can aid in the challenge of assessing indirect discrimination in algorithmic-supported decision-making processes when demographic group data are not available.
@@ 2.2 Bias Detection With and Without Access to Demographic Groups
In the algorithmic fairness literature, identifying systematic differences in the treatment of individuals or groups compared to others is commonly referred to as bias detection. Most bias detection methods assume access to demographic groups such as gender or ethnicity [12, 33, 43, 40, 20, 23]. However, in practice, these labels are often unavailable due to privacy legislation [66]. For instance, the EU’s GDPR restricts the collection of “special categories of data” (article 9) [30], which include data on ethnicity, religion, health, and sexual orientation [^5] Age and gender are not considered special categories of data. See Article 9(1) GDPR. Although, a caveat could exist [67]. The other way around, “political opinions,” “trade union membership,” “genetic,” and “biometric” data are special categories of data but are not protected by EU non-discrimination directives.. The EU AI Act introduces a potential exception to the prohibition on processing special category data: “to the extent that it is strictly necessary for the purpose of ensuring bias detection and correction” [29]. Nonetheless, ongoing legal uncertainties regarding the interplay between the AI Act and the GDPR underscore the need for methods to assess bias in algorithmic systems without relying on demographic data. Algorithms can be biased even in the absence of demographic data, particularly through proxy discrimination, where a set of attributes act as a proxy for the protected attribute [10]. For example, Bayesian Improved Surname Geocoding (BISG) [23] uses Bayes Theorem and data from the United States Census Bureau, including their surname and address, as a proxy for ethnicity or race [10]. However, when information on protected characteristics is not available, the extent to which a characteristic serves as a proxy cannot be investigated and proxy detection methods are therefore not usable in these settings.
There are few alternatives to using data on demographic groups. One set of methods circumvents the problem of limited availability of protected characteristics by using a Rawlsian definition of fairness [59] that maximizes utility for the most disadvantaged group [44, 34, 17]. However, the outcome of these methods is found to not always adhere to more traditional parity-based group fairness metrics [38]. A sec set of alternatives are unsupervised learning methods for fairness [50, 41]. These methods tend to modify traditional clustering algorithms to handle fairness criteria explicitly. Our work contributes to this latter category. The main benefit of this set of methods compared to proxy inferring methods is that they do not require specification of demographic groups beforehand. Thus, unsupervised learning methods are particularly helpful to examine the effects of apparently neutral provisions that might result in indirect discrimination.
@@ 2.3 Auditing Algorithmic Systems
Inspired by established practices in non-algorithmic disciplines, auditing has the potential to mitigate ethical and legal risks in algorithmic systems through both internal and external oversight [58]. Legislators across various jurisdictions are implementing regulations to impose audit frameworks for these systems. For instance, since 2023, New York City has required automated employment decision-making tools to undergo audits by independent third-party auditors [53]. Similarly, since 2024, the EU Digital Services Act mandates annual independent third-party audits for online platforms and search engines with over 45M annual users [28]. In the coming years, the EU AI Act places predominantly self-regulatory obligations on producers and deployers of high-risk AI systems [29].
Implementing effective auditing regimes for algorithmic systems, however, poses practical and cultural challenges. Auditors frequently encounter obstacles such as restricted data access [32], quandaries about which fairness metrics fit the context best [19], and varying opinions on what constitutes a legitimate auditor [32]. Furthermore, auditing reports are not always required to be disclosed, reducing transparency and hindering the development of public knowledge regarding auditing standards. By discussing the audit of DUO’s risk profiling algorithm, we contribute to public knowledge on best practices for auditing algorithmic systems.
3 Unsupervised Bias Detection Tool
This section introduces the theoretical background of the unsupervised bias detection tool. We use examples from our DUO case study to clarify the methodology. We highlight that our research provides a realistic case-study of applying the tool to investigate its usefulness when one lacks protected attribute data. Our study is of both academic and practical interest since the initial algorithm is easy to apply but lacked justification, and because our data is both comprehensive and from a high-stakes policy use-case. Figure 1 provides an overview of all steps in applying the tool.
Figure 1: Schematic overview of the steps involved in applying the unsupervised bias detection tool. The required information is a dataset, a classifier, and a bias metric. Part of the dataset is used to train the Hierarchical Bias-Aware Clustering algorithm [49]. Another part of the dataset is used to test whether differences in the bias metric across clusters are statistically significant.

### v3

@@ Abstract
Algorithms are increasingly used to automate or aid human decisions, yet recent research shows that these algorithms may exhibit bias across legally protected demographic groups. However, data on these groups may be unavailable to organizations or external auditors due to privacy legislation. This paper studies bias detection using an unsupervised bias detection tool when data on demographic groups are unavailable. We collaborated with the Dutch Executive Agency for Education to audit an algorithm that was used to assign risk scores to college students at the national level in the Netherlands between 2012-2023. Our audit covers more than 250,000 students across the country. The unsupervised bias detection tool highlights known disparities between students with a non-European migration background and students with a Dutch or European-migration background. Our contributions are two-fold: (1) we assess bias in a real-world, large-scale, and high-stakes decision-making process by a governmental organization; (2) we provide the unsupervised bias detection tool in an open-source library for others to use to complete bias audits. Our work serves as a starting point for a deliberative assessment by human experts to evaluate potential discrimination in algorithmic decision-making.
@@ 1 Introduction
Between 2012 and 2023, a simple rule-based risk profiling algorithm from the Dutch Executive Agency for Education (DUO) contributed to indirect discrimination in a control process to check whether students were unduly allocated a college grant (Dutch Parliament, 2024b). A rule-based algorithm, consisting of three apparently neutral characteristics of the students (type of education, age, and distance to parents), was used to assign higher risk scores to students enrolled in vocational education, younger students, and students who lived near their parental address. The seemingly impartial risk assessment led to unequal treatment of students with a non-European migration background, remaining unnoticed for over 10 years and affecting more than 350,000 students (Dutch Parliament, 2024a) [^1] Between 2012-2023, more than 350,000 students were assigned a risk score by the algorithm. Our analysis focuses on the student population in two years: 2014 and 2019, which consists of a total of 298,882 students. “Risk” in this context, and throughout the text, refers to the risk of “undue” or improper granting of college grants to students.
@@
Given this, in 2024, the Dutch Minister for Education, Culture and Science apologized on behalf of the Dutch government for indirect discrimination experienced by students with a non-European migration background and announced a EUR 61 million compensation plan (Dutch Parliament, 2024a). This recent discriminatory algorithm scandal in the Netherlands (HEikkilä, 2022) highlights that not only complex statistical methods, such as machine learning, but also simple rule-based algorithms can pose significant risks of embedding discrimination in algorithmic decision-making processes. In a recent interview, the chair of the Dutch Data Protection Authority warned that discriminatory rule-based algorithms are “around every corner” in executive branches of the Dutch government (Rensen, 2024). This type of risk profiling is prohibited by the European Convention of Human Rights (ECHR), the Charter of Fundamental Rights of the European Union (EU) (European Parliament, the Council, and the Commission, 2000), the EU General Data Protection Regulation (GDPR) (European Parliament and the Council, 2016) and the EU AI Act (European Parliament and the Council, 2024), as well as the national Public Administration Law (Dutch Parliament, 1994). Limited institutional enforcement capabilities, along with a lack of awareness, have led to inadequate oversight of algorithmic decision-making processes (Meuwese et al., 2024).
@@
To combat bias in algorithmic systems, the academic community has presented a multitude of ways to define, detect, and mitigate different notions of bias at different points along the algorithmic pipeline (Veale et al., 2018; Barocas et al., 2019; Chouldechova and Roth, 2020; Mehrabi et al., 2021). Most bias mitigation methods require data on demographic groups to prevent discriminatory bias. These demographic groups are typically protected attributes under discrimination law, such as race, gender and age (Ashurst and Weller, 2023). [^2] The exact protected grounds are enshrined in European non-discrimination law, such as the Treaty on the Functioning of the European Union (TFEU), the EU Charter of Fundamental Rights, and specific directives like the Racial Equality Directive (2000/43/EC) and the Employment Equality Directive (2000/78/EC). Note that some provisions do not mention any particular grounds of discrimination, such as Article 20 of the EU Charter of Fundamental Rights which includes openly formulated clauses such as “everyone is equal before the law.” The main protected attributes include: race and ethnicity, gender and sex, religion or belief, disability, age, sexual orientation, nationality, and marital or family status. We will refer to these attributes as demographic groups. However, data on demographic groups are often not available to organizations or external algorithm auditors (Andrus et al., 2021; Centre for Data Ethics and Innovation, 2023). Unsupervised learning can aid in this challenge by detecting groups in the data for which a given notion of equal treatment is violated, avoiding the need for ex-ante specification of demographic or legally protected groups (Misztal-Radecka and Indurkhya, 2021).
@@
In this paper, we use an unsupervised bias detection tool to re-audit the risk profiling algorithm from DUO for bias. The original audit, conducted by Algorithm Audit in 2023, relied on aggregated demographic data on country of birth and country of origin provided by the Dutch national statistics office, Statistics Netherlands (2024b), and is referred to as “supervised bias detection.” For both audits, DUO provided data processed by their algorithm that is not publicly accessible. By combining results from the supervised and unsupervised audit, we assess whether the clusters identified by the unsupervised bias detection tool, which represent potentially unfairly treated groups, correspond to disadvantaged demographic groups resulting from the supervised analysis. This allows for a comparison between an unsupervised learning approach and a supervised approach. We emphasize that the clustering procedure itself does not rely on the available protected demographic data: we only use the aggregated demographic data to illustrate, post hoc, that the flagged clusters align with negatively affected groups. This additional analysis is therefore a validation of the unsupervised clustering methodology, not a replacement for supervised bias detection.
@@
In our context, the bias metric [^3] Throughout our paper, we use the term bias metric to refer to what is commonly called a fairness metric in the algorithmic fairness literature. The bias metric we define reflects the fairness metric of demographic parity (Dwork et al., 2012). requires equal probabilities of “high risk” classification for students with a Dutch or European migration background and those with a non-European migration background. We justify the use of this metric since it assesses the adverse effect of the risk profiling algorithm on different demographic groups: a crucial aspect for investigating indirect discrimination under European non-discrimination law. We find that the cluster with the highest bias returned by the unsupervised bias detection tool is characterized by students who are more often enrolled in vocational education, live close to their parental address, and are relatively more often 15-18 or 25-50 years old. This aligns with demographic groups found to be negatively affected by the risk profiling algorithm, based on aggregated statistics regarding students’ non-European migration background.
Our paper makes two contributions: (1) A large scale audit of indirect discrimination: The DUO audit presents a high-quality and socially relevant dataset, providing insights into the assessment of indirect discrimination in algorithmic decision-making. This study thus contributes to the advancement of public knowledge on auditing algorithmic systems for indirect discrimination. (2) An open-source tool: The unsupervised bias detection method is available as an open-source Python package [^4] https://github.com/NGO-Algorithm-Audit/unsupervised-bias-detection and a web application. [^5] https://algorithmaudit.eu/technical-tools/bdt/#web-app The latter allows non-technical users to apply the tool to their own datasets in order to detect potential bias. The underlying algorithm we provide includes methodological improvements that substantially reduce the false positive rate.
@@
Overall, our paper showcases the potential of unsupervised learning to aid in auditing algorithmic decision-making processes through a real-world use case. We argue that unsupervised bias detection serves as a starting point for a deliberative assessment by human experts to evaluate potential discrimination and unfairness in the process when demographic groups are unavailable.
@@ 2 Literature review
We review a selection of the literature on bias detection and non-discrimination law, focusing on three topics: (1) challenges in detecting and establishing indirect discrimination, (2) bias detection methods with and without access to demographic data, and (3) auditing algorithmic systems for bias.
@@ 2.1 Challenges in detecting and establishing discrimination
Under European non-discrimination law, DUO’s control process is classified as indirect discrimination because a neutral practice disproportionately disadvantaged non-European migrant students (Court of Justice of the European Union, 1986). In algorithms, discrimination often arises from the proxy and correlation challenge (Gerards and Xenidis, 2021). Proxies refer to characteristics that are correlated with a demographic group, leading to disadvantaged outcomes for this group. For example, the Court of Justice of the European Union (CJEU) found that using part-time employment as a basis for differentiation in the context of occupational pension schemes constituted indirect discrimination, since more than 80% of part-time workers in Spain in 2012 were women (Court of Justice of the European Union, 2012). In some cases, courts may accept that differential treatment of a demographic group is acceptable. Under EU law, indirect discrimination can be justified if it is demonstrated that a legitimate aim is being pursued, and the means of achieving that aim are appropriate and necessary (Council of the European Union, 2000a, b, 2004; European Parliament and the Council, 2006). Our paper highlights how unsupervised methods can aid in the challenge of assessing indirect discrimination in algorithmic decision-making processes when demographic group data is not available.
@@ 2.2 Bias detection with and without access to demographic groups
In the algorithmic fairness literature, identifying systematic differences in the treatment of individuals or groups compared to others is commonly referred to as bias detection. Most bias detection methods assume access to demographic groups (Elliot et al., 2009; Hardt et al., 2016; Kusner et al., 2017; Kleinberg et al., 2018; Bellamy et al., 2019; Delaney et al., 2024). However, in practice, these labels are often unavailable due to privacy legislation (Van Bekkum and Borgesius, 2023). For instance, Article 9 in the EU’s GDPR restricts the collection of “special categories of data” (European Parliament and the Council, 2016), which include data on ethnicity, religion, health, and sexual orientation. [^6] Age and gender are not considered special categories of data, see Article 9(1) GDPR. Although, a caveat could exist (van den Brink et al., 2018). The other way around, “political opinions,” “trade union membership,” “genetic,” and “biometric” data are special categories of data but are not protected by EU non-discrimination directives. The EU AI Act introduces a potential exception to the prohibition on processing special category data: “to the extent that it is strictly necessary for the purpose of ensuring bias detection and correction” (European Parliament and the Council, 2024). Nonetheless, ongoing legal uncertainties regarding the interplay between the AI Act and the GDPR underscore the need for methods to assess bias in algorithmic systems without relying on demographic data. Algorithms can be biased even in the absence of demographic data, particularly through proxy discrimination, where a set of attributes acts as a proxy for the protected attribute (Ashurst and Weller, 2023).
Few alternatives exist to approaches that rely on demographic group data. One class of methods circumvents the limited availability of such data by adopting a Rawlsian definition of fairness (Rawls, 2001) that maximizes utility for the most disadvantaged group (Hashimoto et al., 2018; Lahoti et al., 2020; Chai and Wang, 2022). However, the results of these methods do not always adhere to more traditional parity-based group fairness metrics (Islam et al., 2024). A set of alternatives are unsupervised learning methods for fairness (Nasiriani et al., 2019; Krishnakumar et al., 2021). These methods tend to modify traditional clustering algorithms to handle fairness criteria explicitly. Our work contributes to this latter category. Unsupervised learning methods are particularly helpful to examine the effects of apparently neutral provisions that might result in indirect discrimination when demographic groups are not known beforehand.
@@ 2.3 Auditing algorithmic systems
Inspired by established practices in non-algorithmic disciplines, auditing has the potential to mitigate ethical and legal risks in algorithmic systems through both internal and external oversight (Raji et al., 2022). Legislators across various jurisdictions are implementing regulations to impose audit frameworks for these systems. For instance, since 2023, New York City has required automated employment decision-making tools to undergo audits by independent third-party auditors (New York City Council, 2023). Similarly, since 2024, the EU Digital Services Act mandates annual independent third-party audits for online platforms and search engines with more than 45M annual users (European Parliament and the Council, 2022). In the coming years, the EU AI Act places predominantly self-regulatory obligations on producers and deployers of high-risk AI systems (European Parliament and the Council, 2024).
Implementing effective audits for algorithmic systems poses practical and cultural challenges. Auditors frequently encounter obstacles such as restricted data access, differing views on what constitutes a legitimate auditor (Groves et al., 2024), and uncertainty about which fairness metrics best fit the context (Corbett-Davies et al., 2023). Furthermore, auditing reports are not always required to be disclosed, reducing transparency and hindering the development of public knowledge regarding auditing standards. By discussing the audit of DUO’s risk profiling algorithm, we contribute to public knowledge on best practices for auditing algorithmic systems.
3 Insights into the unsupervised bias detection tool
This section introduces the theoretical background of the unsupervised bias detection tool. We use examples from our DUO case study to clarify the methodology. We highlight that our research provides a realistic case-study of applying the tool to investigate its usefulness when one lacks protected attribute data. Our study is of both academic and practical interest since the initial algorithm is easy to apply but lacked justification, and because our data is both comprehensive and from a high-stakes policy use-case. Figure 1 provides an overview of all the steps in applying the tool. We also provide a brief overview of our second contribution, the open source bias detection tool.
Figure 1: Schematic overview of the steps involved in applying the unsupervised bias detection tool. The required information is a dataset, a classifier’s outputs, and a bias metric. Part of the dataset is used to train the Hierarchical Bias-Aware Clustering algorithm (Misztal-Radecka and Indurkhya, 2021). Another part of the dataset is used to test whether differences in the bias metric across clusters are statistically significant.